OpenStudy (ujjwal):

Integrate: This is to be solved by shortest way possible such that it can be solved in a maximum of 1 minute time or 2-3 steps.. The question is mentioned below:

OpenStudy (ujjwal):

\[\int\limits_{?}^{?} dx/\sqrt{(x-1)(x-2)}\]

OpenStudy (ujjwal):

there are no integrating limits..

OpenStudy (experimentx):

you mean indefinite integration right??

OpenStudy (anonymous):

\[ \int\frac{dx}{\sqrt{(x-1)(x-2)}} \]

OpenStudy (experimentx):

try method of completing squares x^2 - 3x +2 = x^2 - 2 * 3/2 x + 9/4 - 9/4 +2

OpenStudy (ujjwal):

yeah @nbouscal that is my qsn

OpenStudy (experimentx):

dont put _{}^{} ... just put \int in you latex for that

OpenStudy (experimentx):

(x-2)(x-1) = (x - 3/2)^2 -1/4 substitute x-3/2 = u it reduces to statndard form of \[ \int \frac{1}{\sqrt{x^ 2 - 1 }} dx\] if you know the formula ... shortcut .. you can solve it in next two steps

OpenStudy (anonymous):

First you want to pull out that 1/4 though, so you'd have \[ \left(\frac{x-3/2}{\sqrt{1/4}}\right)^2-1 \]Right?

OpenStudy (experimentx):

Oops ... sorry .. my mistake!!

OpenStudy (ujjwal):

None of the option contain ln or natural log..

OpenStudy (ujjwal):

so far i know, it leads to natural log or \[\cosh ^{-1}.............something.....\]

OpenStudy (anonymous):

Wolf's solution does use a natural log, and uses basically the same method experimentX suggested.

OpenStudy (ujjwal):

\[a) 2xy (\sqrt{x-1}+\sqrt{x+1}) +c\]\[b) xy(\sqrt{x-1}+\sqrt{x-2})+c\]\[a)2xy(\sqrt{x-1}-\sqrt{x-2})+c\]\[d)\sin ^{-1}\sqrt{(x-1)/(x-2}+c\] thesee are the options.

OpenStudy (ujjwal):

*these

OpenStudy (experimentx):

try differentiating these ... you must get the function inside your integration!!

OpenStudy (experimentx):

try verifying the rest ...

OpenStudy (ujjwal):

I guess none options match..